OpenStudy (anonymous):

Integrate the value and simplify. ∫1/((x-1)√(4x^2-8x+3)) dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ (x-1)\sqrt{4x^2-8x+3} }dx\]

OpenStudy (nincompoop):

Take the integral: integral 1/((x-1) sqrt(4 x^2-8 x+3)) dx For the integrand 1/((x-1) sqrt(4 x^2-8 x+3)), complete the square: = integral 1/((x-1) sqrt((2 x-2)^2-1)) dx For the integrand 1/((x-1) sqrt((2 x-2)^2-1)), substitute u = 2 x-2 and du = 2 dx: = 1/2 integral 1/(sqrt(u^2-1) ((u+2)/2-1)) du For the integrand 1/(sqrt(u^2-1) ((u+2)/2-1)), substitute u = sec(s) and du = tan(s) sec(s) ds. Then sqrt(u^2-1) = sqrt(sec^2(s)-1) = tan(s) and s = sec^(-1)(u): = 1/2 integral (sec(s))/(1/2 (sec(s)+2)-1) ds For the integrand (sec(s))/(1/2 (sec(s)+2)-1), do long division: = 1/2 integral 2 ds The integral of 2 is 2 s: = s+constant Substitute back for s = sec^(-1)(u): = sec^(-1)(u)+constant Substitute back for u = 2 x-2: = sec^(-1)(2 x-2)+constant Which is equivalent for restricted x values to: Answer: -tan^(-1)(1/sqrt(4 x^2-8 x+3))+constant

OpenStudy (kirbykirby):

ok well you beat me to it lol

OpenStudy (anonymous):

You two are too kind.